OpenStudy (anonymous):

Can someone please check my answers ?? i will fan and medal(: Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 + 2x + y2 + 4y = 20

OpenStudy (anonymous):

(X-h)^2 + (y - K)^2 = r^2 (x^2 + 2x + 1) + (y^2 + 2(2y) +4) = 20 + 1 + 4 (perfect squares) Next step is easy: you compare (x+1)^2 + (y+2)^2 = 5^2 with the standard form of the circle with centre (h,k) and radius r (x-h)^2+(y-k)^2 = r^2 and decide what the values of h, k and r are. From (y+2) = y-k, subtract y from both sides, so y-y+2 = y-y-k or 2=-k, multiply by -1 to get -2=k, so we have h=-1, k=-2, and r=5 From the standard equation of a circle, (x-h)^2+(y-k)^2=r^2 we therefore found center = (h,k)=-1,-2, and radius = r =5

OpenStudy (anonymous):

@zepdrix @jim_thompson5910

jimthompson5910 (jim_thompson5910):

I think you meant to say center = (-1,-2)

jimthompson5910 (jim_thompson5910):

if so, you have the correct answer

OpenStudy (anonymous):

yeah i did my bad , but is everything in the right order and everything ? and seem correct @jim_thompson5910

jimthompson5910 (jim_thompson5910):

everything else looks perfect. Nice work

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Thank you!